OpenStudy (anonymous):

multiply: (x+2)(x+4)= i came up with the answer of: 2x+4+8 is this correct

OpenStudy (anonymous):

No, this is not correct

OpenStudy (anonymous):

x^2+6x+8

OpenStudy (lone001):

=x^2+6*x+8

OpenStudy (anonymous):

Always do: first x first + first x second + second x second x^2 + 4x + 2x + 8 x^2 + 6x + 8

OpenStudy (anonymous):

x*x+4*x+2*x+4*2

OpenStudy (anonymous):

no do what hickninga does

OpenStudy (anonymous):

I meant first x first + first x second + second x first + second x second

OpenStudy (anonymous):

imran wrote the exact same thing as hickninja

OpenStudy (amistre64):

just multiply the things like normal multiplication ..... which is what multiplyind IS :)

OpenStudy (anonymous):

that, too.

OpenStudy (anonymous):

(a + b)(c + d) = a(c + d) + b(c + d) = ac + ad + bc + bd. (or first x first, first x second...)

OpenStudy (amistre64):

x +2 x +4 ------- 4x +8 x^2 +2x ------------ x^2 +6x +8

OpenStudy (anonymous):

amistre64 thank you for the clarification

OpenStudy (anonymous):

i gave medalsss

OpenStudy (amistre64):

youre welcome :)